six times Paul's age plus three times Annette's age equals 72. Annette's age is also four times Paul's age. How old is Paul?
Answer by JulietG(1812)

6P + 3A = 72
A = 4P
Substitute the known value of A from the second equation into the first.
6P + 3(4P) = 72
6P + 12P = 72
18P = 72
Divide each side by 18
P = 4
.
If Paul is 4, then Annette is 16.